# Generalised rank-constrained approximations of Hilbert–Schmidt operators on separable Hilbert spaces and applications

### v1

@@ 5 Bounded solutions and approximations
Theorem 3.2 prescribes a set of solutions that are of the form (11). As shown in Section 4, this set can contain unbounded solutions. It is even possible that some solution of the form (11) are bounded and others are unbounded. It is natural to ask when all solutions are bounded. In Section 5.1 we give necessary and sufficient conditions for (11) to have only bounded solutions. In the case they are not bounded, we use the concept of outer inverses to construct suitable bounded approximations in Section 5.2.
@@ 5.1 Necessary and sufficient conditions for bounded solutions
We recall that Proposition A.5 item (iv) gives a sufficient condition for (11) to consist of only bounded solutions, namely that $\operatorname{ran}{C}$ is closed in $\mathcal{H}_{2}$. An even stronger condition is given next.
@@
    $\ker{\hat{X}}$ is closed in $\operatorname{dom}{C^{\dagger}}$,
@@
    $C(\ker{Z}\cap\ker{C}^{\perp})$ is closed in $\operatorname{ran}{C}$,
@@
The equivalence of items (i) and (ii) is proven similarly as it is proven for linear functionals, i.e. when $r=1$, that continuity is equivalent to having a closed kernel. We outline a proof in the appendix as Lemma B.1. To show the equivalence between items (ii) and (iii) we compute $\ker{\hat{X}}$. As $\operatorname{ran}{C^{\dagger}}=\ker{C}^{\perp}$, c.f. Proposition A.5 item (i),
@@
We give a final sufficient condition, a lower bound condition, that is weaker than $C$ having closed range. Since $C$ is bounded, $\ker{C}$ is closed. We recall that the quotient space $\mathcal{H}_{1}/\ker{C}$ then is a Banach space with norm $\lVert h+\ker{C}\rVert=\inf\{\lVert h+c\rVert:\ c\in\ker{C}\}$. The quotient map $Q\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{1}/\ker{C})$, $h\mapsto h+\ker{C}$ is continuous. See for example [16, Section III.4].
@@
Let us denote $W\coloneqq\ker{Z}\cap\ker{C}^{\perp}$. We show that $C(W)$ is closed in $\mathcal{H}_{4}$. Then $C(W)$ is also closed in $\operatorname{ran}{C}$ and we may apply Proposition 5.2. Now, we can decompose $C=\tilde{C}\circ Q$ for some $\tilde{C}\in\mathcal{B}(\mathcal{H}_{1}/\ker{C},\mathcal{H}_{2})$. Let $(w_{n})_{n}$ be a sequence in $W$ such that $(Cw_{n})_{n}=(\tilde{C}(w_{n}+\ker{C}))_{n}$ converges to $k\in\mathcal{H}_{4}$. We need to show $k=Cw$ for some $w\in W$. By assumption, $\lVert w_{n}-w_{m}+\ker{C}\rVert\leq c^{-1}\lVert C(w_{n}-w_{m})\rVert$, so $(w_{n}+\ker{C})_{n}$ is a Cauchy sequence in $\mathcal{H}_{1}/\ker{C}$. But $\mathcal{H}_{1}/\ker{C}$ is complete, so $w_{n}+\ker{C}$ converges to some $h+\ker{C}\in\mathcal{H}_{1}/\ker{C}$. As $W$ and $\ker{C}$ are orthogonal, their direct sum $W+\ker{C}$ is closed, c.f. [4, Section 9.2.(A)]. Hence $h+\ker{C}\in W+\ker{C}$, i.e. $h+\ker{C}=w+\ker{C}$ for some $w\in W$. Now, $Cw=Ch=\tilde{C}(h+\ker{C})=\lim_{n}\tilde{C}(w_{n}+\ker{C})=k$, by continuity of $\tilde{C}$. ∎
In case $C$ has closed range, the operator $\tilde{C}$ in the proof of Corollary 5.3 also has closed range. But $\tilde{C}$ is also injective, so it has a bounded inverse on its range, by the inverse mapping theorem, c.f. [16, Theorem III.12.5]. Hence
@@ 5.2 Bounded approximations
$\inf\{\lVert M-BXC\rVert_{HS}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\},$ (20)
@@
As $\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$, this infimum is bounded from below by the minimum in (10). In the case that $\operatorname{ran}{C}$ is closed, the problems (10)-(12) and (20)-(21) coincide by Remark 3.5. Let us therefore assume that $\operatorname{ran}{C}$ is not closed, so $\ker{C}^{\perp}$ is infinite-dimensional by Lemma 5.1. We now show that in this case the infimum (20) is not attained in general, but is still equal to the minimum in (10). To do so, we construct a minimizing sequence $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ to (20)-(21) such that $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$. We do this under the assumption that $\mathcal{Y}_{r}\not=\emptyset$ and hence that a solution $\hat{X}$ of the form (11) actually exists.
Given $\hat{X}$ of the form (11), we construct the approximating sequence $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ based on the approximation of the unbounded Moore–Penrose inverse $C^{\dagger}$ by outer inverses as proposed by [25]. We briefly repeat the construction and its properties here, but refer to [25] for elaboration.
Let $(Y_{n})_{n}$ be an increasing sequence of finite-dimensional subspaces of $\overline{\operatorname{ran}{C}}$ such that $\overline{\cup_{n}Y_{n}}=\overline{\operatorname{ran}{C}}$. We define $X_{n}\coloneqq C^{*}Y_{n}$ and let $P_{n}$, $Q_{n}$ be the orthogonal projectors onto $Y_{n}$ and $X_{n}$ respectively. We then set $C_{n}\coloneqq P_{n}C$, so $C_{n}$ has finite rank. It can be shown that $(X_{n})_{n}$ is an increasing sequence of finite-dimensional subspaces, that $\overline{\cup_{n}X_{n}}=\ker{C}^{\perp}$ and that $\ker{C_{n}}=X_{n}$, $\operatorname{ran}{C_{n}}=Y_{n}$ so that ${\left.\kern-1.2ptC_{n}\vphantom{\big{|}}\right|_{X_{n}}}:X_{n}\rightarrow Y_{n}$ is a boundedly invertible operator. Hence, we can define the finite-rank operator $C_{n}^{\#}:\mathcal{H}_{2}\rightarrow\mathcal{H}_{1}$ as,
@@
The sequence $(C_{n})_{n}\subset\mathcal{B}_{00}(\mathcal{H}_{2},\mathcal{H}_{1})$ satisfies,
@@
Let us now, under the condition $\mathcal{Y}_{r}\not=\emptyset$, set
@@
In 3.1 suppose $\mathcal{Y}_{r}\not=\emptyset$ and let $\hat{X}$ be given by (11). With $\hat{X}_{n}$ defined in (22), it holds that $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$, $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$ and $B\hat{X}_{n}C\rightarrow B\hat{X}C$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Furthermore, $\hat{X}_{n}$ satisfies (21).
In Section 4 it is shown that it may happen that all solutions $\hat{X}$ of (10) are unbounded. Since the infimum in (20) and the minimum in (10) are equal by Proposition 5.5, it then follows that the infimum in (20) is not attained in this case. Indeed, if it were attained by some bounded $\tilde{X}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$, then $\tilde{X}$ would be a bounded minimiser of (10), contradicting the fact that in this case all solutions are unbounded.
@@
Let us first notice that, as in the proof of Lemma 3.7 item (i), we can use the first equation of (14) and the line preceding it to find that
$B\hat{X}C=(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}P_{\ker{C}^{\perp}},$
and we can use Lemma A.1 to find that $B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$ is bounded. Now, being a composition of bounded operators, $\hat{X}_{n}$ is bounded. Also, $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$ since $C_{n}^{\#}\rightarrow C^{\dagger}$ pointwise on $\operatorname{dom}{C^{\dagger}}=\operatorname{dom}{\hat{X}}$ by Theorem 5.4.
As the $(X_{n})_{n}$ are increasing, we can take an ONB $(e_{i})_{i}$ of $\ker{C}^{\perp}$ such that for all $n$, $\{e_{1},\ldots,e_{n}\}$ is an ONB of $X_{n}$ and $e_{i}\in X_{n}^{\perp}$, $i>n$. Now, $Q_{n}P_{\ker{C}^{\perp}}=Q_{n}$ since $\ker{Q_{n}}^{\perp}=X_{n}\subset\overline{\operatorname{ran}{C}}$. By Theorem 5.4, $C_{n}^{\#}=Q_{n}C^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$, so with (36), $C_{n}^{\#}C=Q_{n}C^{\dagger}C=Q_{n}$, and using also (34),(36),(37) it follows that,
$B\hat{X}_{n}C$ $=BB^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C_{n}^{\#}C=(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}Q_{n},$
$P_{\ker{B}^{\perp}}\hat{X}_{n}P_{\overline{\operatorname{ran}{C}}}$ $=B^{\dagger}B\hat{X}_{n}CC^{\dagger}=B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}Q_{n}C^{\dagger}=\hat{X}_{n}\quad\text{on }\operatorname{dom}{C^{\dagger}}.$
The second equation shows that (21) holds, as $\operatorname{dom}{C^{\dagger}}\subset\mathcal{H}_{2}$ densely and $\hat{X}_{n}$ is continuous, while with the first we compute,
$\lVert B\hat{X}C-B\hat{X}_{n}C\rVert_{HS}^{2}$ $=\lVert(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}(P_{\ker{C}^{\perp}}-Q_{n})\rVert_{HS}^{2}$
@@
One could also start with an increasing sequence $(X_{n})_{n}$ and define $Y_{n}=CX_{n}$ instead of choosing $(Y_{n})_{n}$ and setting $X_{n}=C^{*}Y_{n}$. So $\operatorname{ran}{P_{n}}=C(\operatorname{ran}{Q}_{n})$ and we could try to construct the bounded approximations $\tilde{C}_{n}\coloneqq C^{\dagger}P_{n}$ to $C^{\dagger}$. If $\overline{\cup_{n}X_{n}}=\ker{C}^{\perp}$, then still $\overline{\cup_{n}Y_{n}}=\overline{\operatorname{ran}{C}}$ and the pointwise convergence $\tilde{C}_{n}\rightarrow C^{\dagger}$ holds on $\operatorname{dom}{C^{\dagger}}$. But $C^{\dagger}P_{n}Ce_{i}$ is not equal to 0 for $i>n$, so $\tilde{C}_{n}C=C^{\dagger}P_{n}C$ is no longer equal to $Q_{n}$, which is what we used to show the $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ convergence in Proposition 5.5. In fact, the sequence,
@@
need not be converging in $n$. So the statement on $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ convergence will not hold in general for this alternative approximation procedure.
@@
The approximate minimisers $\hat{X}^{(m)}$ constructed in Section 3.2 may be unbounded. The procedure outlined before Proposition 5.5 can be used for each $m$ to approximate $\hat{X}^{(m)}$ with bounded approximations $\hat{X}^{(m)}_{n}$. Then $B\hat{X}^{(m)}_{n}C\rightarrow B\hat{X}^{(m)}C$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ as $n\rightarrow\infty.$ In particular $B\hat{X}^{(m)}_{m}C\rightarrow Y$ for some rank-$r$ reduced SVD of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$. Thus $\hat{X}^{(m)}_{m}$ is a sequence of bounded approximate minimisers.
@@
$\min\{\lVert M^{*}-C^{*}XB^{*}\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}(C^{*},B^{*})\},$ (23)
@@
$\min\{\lVert M-BXC\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}(B,C)\}$ $=\inf\{\lVert M-BXC\rVert_{HS}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}$
$=\inf\{\lVert M^{*}-C^{*}X^{*}B^{*}\rVert_{HS}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}$
$=\inf\{\lVert M^{*}-C^{*}XB^{*}\rVert_{HS}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})\}$
$=\min\{\lVert M^{*}-C^{*}XB^{*}\rVert_{HS}^{2}:\ X\in\mathcal{X}_{r}(C^{*},B^{*})\}.$
Here we used that $(BXC)^{*}=C^{*}X^{*}B^{*}$ holds true when $B,X,C$ are bounded. When $\mathcal{Y}_{r}(B,C)=\emptyset$, the minimum of the left hand side becomes an infimum that is not attained. It is, however, still possible that $\mathcal{Y}_{r}(C^{*},B^{*})\not=\emptyset$. In this case the adjoint problem (23) still has a solution and allows us to recover the minimal Hilbert–Schmidt approximation error, even though (10) does not admit a minimal solution. Another reason for why (23) can be useful is that unlike (10), which has only unbounded solutions, (23) admits a bounded solution. In this case, the adjoint problem allows to recover the minimal Hilbert–Schmidt approximation error using a bounded operator. Section 6 shows an example of the latter case.

### v3

@@ 5 Bounded solutions and approximations
Theorem 3.2 prescribes a set of solutions that are of the form (11). As shown in Section 4, this set can contain unbounded solutions. It is even possible that some solutions of the form (11) are bounded and others are unbounded. It is natural to ask when all solutions are bounded. In Section 5.1 we give necessary and sufficient conditions for (11) to have only bounded solutions. In the case where they are not bounded, we use the concept of outer inverses to construct suitable bounded approximations in Section 5.2.
@@ 5.1 Necessary and sufficient conditions for bounded solutions
We recall that Remark 3.5(i) gives a sufficient condition for (11) to consist of only bounded solutions, namely that $\operatorname{ran}{C}$ is closed in $\mathcal{H}_{2}$. An even stronger condition is given next.
@@
    $\ker{\hat{X}}$ is closed in $\operatorname{dom}{C^{\dagger}}$.
@@
    $C(\ker{Z}\cap\ker{C}^{\perp})$ is closed in $\operatorname{ran}{C}$.
@@
The equivalence of items (i) and (ii) is proven similarly as one proves the equivalence of continuity and having closed range for linear functionals, i.e. when $r=1$. We outline a proof in the appendix as Lemma B.1. To show the equivalence between items (ii) and (iii) we compute $\ker{\hat{X}}$. As $\operatorname{ran}{C^{\dagger}}=\ker{C}^{\perp}$, c.f. Proposition A.5 item (i), we have
@@
We give a final sufficient condition, namely a lower bound condition, that is weaker than $C$ having closed range. Since $C$ is bounded, $\ker{C}$ is closed. We recall that the quotient space $\mathcal{H}_{1}/\ker{C}$ then is a Banach space with norm $\lVert h+\ker{C}\rVert=\inf\{\lVert h+c\rVert:\ c\in\ker{C}\}$, and that the quotient map $Q\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{1}/\ker{C})$, $h\mapsto h+\ker{C}$ is continuous. See for example [15, Section III.4].
@@
Let us denote $W\coloneqq\ker{Z}\cap\ker{C}^{\perp}$. We show that $C(W)$ is closed in $\mathcal{H}_{2}$. Then $C(W)$ is also closed in $\operatorname{ran}{C}$ and we may apply Proposition 5.2. Now, we can decompose $C=\tilde{C}\circ Q$ for some $\tilde{C}\in\mathcal{B}(\mathcal{H}_{1}/\ker{C},\mathcal{H}_{2})$. Let $(w_{n})_{n}$ be a sequence in $W$ such that $(Cw_{n})_{n}=(\tilde{C}(w_{n}+\ker{C}))_{n}$ converges to $k\in\mathcal{H}_{4}$. We need to show $k=Cw$ for some $w\in W$. By assumption, $\lVert w_{n}-w_{m}+\ker{C}\rVert\leq c^{-1}\lVert C(w_{n}-w_{m})\rVert$, so $(w_{n}+\ker{C})_{n}$ is a Cauchy sequence in $\mathcal{H}_{1}/\ker{C}$. But $\mathcal{H}_{1}/\ker{C}$ is complete, so $w_{n}+\ker{C}$ converges to some $h+\ker{C}\in\mathcal{H}_{1}/\ker{C}$. As $W$ and $\ker{C}$ are orthogonal, their direct sum $W+\ker{C}$ is closed, c.f. [4, Section 9.2.(A)]. Hence $h+\ker{C}\in W+\ker{C}$, i.e. $h+\ker{C}=w+\ker{C}$ for some $w\in W$. Now, $Cw=Ch=\tilde{C}(h+\ker{C})=\lim_{n}\tilde{C}(w_{n}+\ker{C})=k$, by continuity of $\tilde{C}$. ∎
In case $C$ has closed range, the operator $\tilde{C}$ in the proof of Corollary 5.3 also has closed range. But $\tilde{C}$ is also injective, so it has a bounded inverse on its range, by the inverse mapping theorem, c.f. [15, Theorem III.12.5]. Hence
@@ 5.2 Bounded approximations
$\inf\{\lVert M-BXC\rVert_{L_{2}}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\},$ (20)
@@
As $\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$, this infimum is bounded from below by the minimum in (10). In the case that $\operatorname{ran}{C}$ is closed, the problems (10)-(12) and (20)-(21) coincide by Remark 3.5(i). Let us therefore assume that $\operatorname{ran}{C}$ is not closed, so $\ker{C}^{\perp}$ is infinite-dimensional by Lemma 5.1. We now show that in this case the infimum (20) is not attained in general, but is still equal to the minimum in (10). To do so, we construct a minimizing sequence $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ to (20)-(21) such that $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$. We do this under the assumption that $\mathcal{Y}_{r}\not=\emptyset$ and hence that a solution $\hat{X}$ of the form (11) exists.
Given $\hat{X}$ of the form (11), we construct the approximating sequence $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$ based on the approximation of the unbounded Moore–Penrose inverse $C^{\dagger}$ by outer inverses as proposed by [25]. We briefly repeat the construction and its properties here, but refer to [25] for elaboration, see also [18, Theorem 3.24].
Let $(Y_{n})_{n}$ be an increasing sequence of finite-dimensional subspaces of $\overline{\operatorname{ran}{C}}$ such that $\overline{\cup_{n}Y_{n}}=\overline{\operatorname{ran}{C}}$. We define $X_{n}\coloneqq C^{*}Y_{n}$ and let $P_{n}$, $Q_{n}$ be the orthogonal projectors onto $Y_{n}$ and $X_{n}$ respectively. We then set $C_{n}\coloneqq P_{n}C$, so $C_{n}$ has finite rank. It can be shown, c.f. [25, p114], that $(X_{n})_{n}$ is an increasing sequence of finite-dimensional subspaces, that $\overline{\cup_{n}X_{n}}=\ker{C}^{\perp}$ and that $\ker{C_{n}}^{\perp}=X_{n}$, $\operatorname{ran}{C_{n}}=Y_{n}$ so that ${\left.\kern-1.2ptC_{n}\vphantom{\big{|}}\right|_{X_{n}}}:X_{n}\rightarrow Y_{n}$ is a boundedly invertible operator. Hence, we can define the finite-rank operator $C_{n}^{\#}:\mathcal{H}_{2}\rightarrow\mathcal{H}_{1}$ as
@@
The sequence $(C_{n})_{n}\subset\mathcal{B}_{00}(\mathcal{H}_{2},\mathcal{H}_{1})$ satisfies the following:
@@
Under the condition $\mathcal{Y}_{r}\not=\emptyset$, let us now set
@@
In 3.1 suppose $\mathcal{Y}_{r}\not=\emptyset$ and let $\hat{X}$ be given by (11). With $\hat{X}_{n}$ defined in (22), it holds that $\hat{X}_{n}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})$, $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$ and $B\hat{X}_{n}C\rightarrow B\hat{X}C$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Furthermore, $\hat{X}_{n}$ satisfies (21). It follows that the infimum in (20) and minimum in (10) agree.
In Section 4 it is shown that in some cases, all solutions $\hat{X}$ of (10) can be unbounded. Since the infimum in (20) and the minimum in (10) are equal by Proposition 5.5, it then follows that the infimum in (20) is not attained in this case. Indeed, if it were attained by some bounded $\tilde{X}\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\subset\mathcal{X}_{r}$, then $\tilde{X}$ would be a bounded minimiser of (10), contradicting the fact that in this case all solutions are unbounded.
@@
Let us define $Y\coloneqq(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})$. As in the proof of Lemma 3.7 item (i), we can use the first equation of (14) and the line preceding it to find that
$B\hat{X}C=B(B^{\dagger}YC^{\dagger})C=YC^{\dagger}C=YP_{\ker{C}^{\perp}}=(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}P_{\ker{C}^{\perp}},$
and we can use Lemma A.1 to find that $B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}$ is bounded. Now, being a composition of bounded operators, $\hat{X}_{n}$ is bounded. Also, $\hat{X}_{n}\rightarrow\hat{X}$ pointwise on $\operatorname{dom}{\hat{X}}$ since $C_{n}^{\#}\rightarrow C^{\dagger}$ pointwise on $\operatorname{dom}{C^{\dagger}}=\operatorname{dom}{\hat{X}}$ by Theorem 5.4(ii).
Now, $Q_{n}P_{\ker{C}^{\perp}}=Q_{n}$ since $\ker{Q_{n}}^{\perp}=X_{n}\subset\overline{\operatorname{ran}{C}}$. By Theorem 5.4, $C_{n}^{\#}=Q_{n}C^{\dagger}$ on $\operatorname{dom}{C^{\dagger}}$. Combining this with (38) yields $C_{n}^{\#}C=Q_{n}C^{\dagger}C=Q_{n}$. Using also (36), (38), and (39), it follows that
$B\hat{X}_{n}C$ $=BB^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}C_{n}^{\#}C=(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}Q_{n},$ (23)
$P_{\ker{B}^{\perp}}\hat{X}_{n}P_{\overline{\operatorname{ran}{C}}}$ $=B^{\dagger}B\hat{X}_{n}CC^{\dagger}=B^{\dagger}(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}Q_{n}C^{\dagger}=\hat{X}_{n}\quad\text{on }\operatorname{dom}{C^{\dagger}}.$ (24)
Equation (24) shows that (21) holds, as $\operatorname{dom}{C^{\dagger}}\subset\mathcal{H}_{2}$ densely and $\hat{X}_{n}$ is continuous. As the $(X_{n})_{n}$ are increasing, we can take an ONB $(e_{i})_{i}$ of $\ker{C}^{\perp}$ such that for all $n$, $\{e_{1},\ldots,e_{n}\}$ is an ONB of $X_{n}$ and $e_{i}\in X_{n}^{\perp}$, $i>n$. Thus, we compute with (23),
$\lVert B\hat{X}C-B\hat{X}_{n}C\rVert_{L_{2}}^{2}$ $=\lVert(P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}})_{r}(P_{\ker{C}^{\perp}}-Q_{n})\rVert_{L_{2}}^{2}$
@@
One could also start with an increasing sequence $(X_{n})_{n}$ and define $Y_{n}=CX_{n}$ instead of choosing $(Y_{n})_{n}$ and setting $X_{n}=C^{*}Y_{n}$. In this case the orthogonal projector $P_{n}$ onto $Y_{n}$ satisfies $\operatorname{ran}{P_{n}}=C(\operatorname{ran}{Q}_{n})$ and we could try to construct bounded approximations $\tilde{C}_{n}\coloneqq C^{\dagger}P_{n}$ that converge pointwise to $C^{\dagger}$. If $\overline{\cup_{n}X_{n}}=\ker{C}^{\perp}$, then still $\overline{\cup_{n}Y_{n}}=\overline{\operatorname{ran}{C}}$ and the pointwise convergence $\tilde{C}_{n}\rightarrow C^{\dagger}$ holds on $\operatorname{dom}{C^{\dagger}}$. But $C^{\dagger}P_{n}Ce_{i}$ is not equal to 0 for $i>n$, so $\tilde{C}_{n}C=C^{\dagger}P_{n}C$ is no longer equal to $Q_{n}$, which is what we used to show the $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ convergence statement in Proposition 5.5. In fact, the sequence,
@@
need not converge as $n\to\infty$. Thus the $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ convergence statement in Proposition 5.5 will not hold in general for this alternative approximation procedure.
@@
The approximate minimisers $(\hat{X}^{(m)})_{m}$ of some SVD $Y$ of $P_{\overline{\operatorname{ran}{B}}}MP_{\ker{C}^{\perp}}$ constructed in Section 3.2 may be unbounded. The procedure outlined before Proposition 5.5 can be used for each $m$ to approximate $\hat{X}^{(m)}$ with bounded approximations $(\hat{X}^{(m)}_{n})_{n}$. Then $B\hat{X}^{(m)}_{n}C\rightarrow B\hat{X}^{(m)}C$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$ as $n\rightarrow\infty.$ In particular, for each $m\in\mathbb{N}$ there exists $n_{m}\in\mathbb{N}$ such that $\lVert B\hat{X}^{(m)}_{n}C-B\hat{X}^{(m)}C\rVert_{L_{2}}<m^{-1}$ for all $n\geq n_{m}$. Hence $B\hat{X}^{(m)}_{n_{m}}C\rightarrow Y$ in $L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Thus $(\hat{X}^{(m)}_{n_{m}})_{m}$ is a sequence of bounded approximate minimisers. Since $B(P_{\ker{B}^{\perp}}\hat{X}_{n_{m}}^{(m)}P_{\overline{\operatorname{ran}{C}}})C=B\hat{X}_{n_{m}}^{(m)}C$ by (36), (38) and (39), and since $P_{\ker{B}^{\perp}}\hat{X}_{n_{m}}^{(m)}P_{\overline{\operatorname{ran}{C}}}$ satisfies (21), we can thus construct a sequence of bounded minimisers which also satisfy the minimality property (21).
@@
$\min\{\lVert M^{*}-C^{*}XB^{*}\rVert_{L_{2}}:\ X\in\mathcal{X}_{r}(C^{*},B^{*})\},$ (25)
@@
$\min\{\lVert M-BXC\rVert_{L_{2}}^{2}:\ X\in\mathcal{X}_{r}(B,C)\}$ $=\inf\{\lVert M-BXC\rVert_{L_{2}}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}$
$=\inf\{\lVert M^{*}-C^{*}X^{*}B^{*}\rVert_{L_{2}}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{2},\mathcal{H}_{3})\}$
$=\inf\{\lVert M^{*}-C^{*}XB^{*}\rVert_{L_{2}}^{2}:\ X\in\mathcal{B}_{00,r}(\mathcal{H}_{3},\mathcal{H}_{2})\}$
$=\min\{\lVert M^{*}-C^{*}XB^{*}\rVert_{L_{2}}^{2}:\ X\in\mathcal{X}_{r}(C^{*},B^{*})\}.$
Here we used that $(BXC)^{*}=C^{*}X^{*}B^{*}$ holds true when $B,X,C$ are bounded. When $\mathcal{Y}_{r}(B,C)=\emptyset$, the minimum of the left hand side becomes an infimum that is not attained. It is, however, still possible that $\mathcal{Y}_{r}(C^{*},B^{*})\not=\emptyset$. In this case the adjoint problem (25) still has a solution and allows us to recover the minimal Hilbert–Schmidt approximation error, even though (10) does not admit a minimal solution. Another case in which (25) can be useful is when (10) only has unbounded solutions but (25) admits a bounded solution. In this case, the adjoint problem allows to attain the minimal Hilbert–Schmidt approximation error using a bounded operator. Section 6 shows an example of the latter case.